2. Tantalum is a rare, hard, blue-gray, lustrous transition metal that is highly corrosion-resistant. It has a density of 16.4 grams per cubic centimeter and an atomic weight of 180.948 g/mol. What is the atomic radius of Ta if it adopts the body centered cubic unit structure? Express your answer in Å.
1
Expert's answer
2021-09-07T23:28:55-0400

"M= 180.948 \;g/mol \\

ρ= 16.4 \;g/cm^3"

N=2 for BCC

For BCC

"a= \frac{4r}{\sqrt{3}} \\

ρ = \frac{N \times M}{6.023 \times 10^{23} \times a^3} \\

r= (\frac{ 2 \times 180.948}{16.4 \times 6.023 \times 10^{23}})^{1/3} \\

= 3.3212 \times 10^{-8} \;cm \\

= 3.3212 \times 10^{-10} \;m \\

= 3.32 \;Å"